Write an equation of the line through the given points.

(2, -4),(0, -3)

Any help with the given equation would Greatly appreciated.

Answer by checkley71(8403)

Find the slope
(y2-y1)/(x2-x1)=(-3+4)/(0-2)=1/-2=-1/2 Answer for the slope.
-3=-1/2*0+b
-3=0+b
b=-3 thus the equation is:
y=-x/2-3 answer.